# Asymptotic Stability of 3D Out-flowing Compressible Viscous Fluid under Non-Spherical Perturbation

### v1

@@ Abstract
We study an outflow problem for the $3$-dimensional isentropic compressible Navier-Stokes equations. The fluid under consideration occupies the exterior domain of unit ball, $\Omega=\{x\in\mathbb{R}^{3}\,|\,|x|\geq 1\}$, and it is flowing out from $\Omega$ at a constant speed $|u_{b}|$, in the normal direction to the boundary surface $\partial\Omega$. It is shown in Hashimoto-Matsumura(2021) that if the fluid velocity at the far-field is assumed to be zero, and $|u_{b}|$ is sufficiently small, then there exists a unique spherically symmetric stationary solution $(\tilde{\rho},\tilde{u})(x)$. In this paper, we prove that $(\tilde{\rho},\tilde{u})(x)$ is asymptotically stable in time, under small, possibly non-spherically symmetric initial perturbations.
@@ 1 Introduction
$\rho\{\partial_{t}u+(u\cdot\nabla)u\}+\nabla P(\rho)=\mu\Delta u+(\mu+\lambda)\nabla\textnormal{{div}}u\equiv Lu.$ (1.1b)
We study the asymptotic behaviour of a solution $(\rho,u)$ to (1.1) in an unbounded exterior domain $\Omega\vcentcolon=\{x=(x_{1},x_{2},x_{3})^{\top}\in\mathbb{R}^{3}\;|\;|x|>1\}$. Here $\rho>0$ is the mass density; $u=(u^{1},u^{2},u^{3})^{\top}$ is the velocity of gas; $P(\rho)=K\rho^{\gamma}\;(K>0,\gamma\geq 1)$ is the pressure with the adiabatic exponent $\gamma$; $\mu$ and $\lambda$ are constant called viscosity-coefficients satisfying $\mu>0$ and $2\mu+3\lambda\geq 0$.
In this paper, we consider the model where fluid is flowing out at a constant velocity from a solid inner sphere centred at the origin of radius $|x|=1$. Its corresponding boundary condition is given as:
[…]
For equations (1.1)–(1.2), we prescribe the initial data $(\rho_{0},u_{0})(x)$ such that
$\rho(x,0)=\rho_{0}(x)>0,\quad u(x,0)=u_{0}(x),$ (1.3a)
$\lim_{|x|\to\infty}(\rho_{0}(x),u_{0}(x))=(\rho_{+},0),\quad 0<\rho_{+}<\infty.$ (1.3b)
In addition, (1.3a) is compatible with the boundary data (1.2a) which states that,
@@
Since the characteristic speed of (1.1a) is negative on the boundary due to (1.2a), no boundary condition for $\rho$ at $x\in\partial\Omega$ is necessary for the well-posedness of the initial boundary value problem (1.1)–(1.4).
Spehrically symmetric steady-state solution
Consider instead the general multi-dimensional exterior domain $\Omega=\{x\in\mathbb{R}^{n}\,|\,|x|\geq 1\}$ where $n\geq 2$. Let $(\tilde{\rho},\tilde{u})$ be the steady state solution of (1.1)–(1.4), solving the system:
$\textnormal{{div}}\big{(}\tilde{\rho}\tilde{u}\big{)}=0$ for $\text{for }\ x\in\Omega,$ (1.5a)
$\tilde{\rho}(\tilde{u}\cdot\nabla)\tilde{u}+\nabla P(\tilde{\rho})=L\tilde{u}$ for $\text{for }\ x\in\Omega,$ (1.5b)
$\tilde{u}(x)=u_{b}\frac{x}{|x|}$ for $\text{for }\ x\in\partial\Omega,$ (1.5c)
$(\tilde{\rho},\tilde{u})(x)\to(\rho_{+},0)$ as |$\text{as }\ |x|\to\infty.$ (1.5d)
Since the boundary and far-field conditions (1.5c)–(1.5d) are spherically symmetric, if the unique solution $(\tilde{\rho},\tilde{u})$ to (1.5) exists, then $(\tilde{\rho},\tilde{u})$ must also be spherically symmetric. This means that $(\tilde{\rho},\tilde{u})$ can be expressed in the form:
$\tilde{\rho}(x)=\tilde{\rho}(r),\quad\tilde{u}(x)=\tilde{U}(r)\frac{x}{r},\qquad\text{where }\ r\equiv|x|\geq 1,$ (1.6)
for some function $\tilde{U}(r)\vcentcolon[1,\infty)\to\mathbb{R}$. Then the system (1.5) is reformulated as
$\partial_{r}\big{(}r^{n-1}\tilde{\rho}\tilde{U}\big{)}=0$ for $\text{for }\ r\geq 1,$ (1.7a)
$\tilde{\rho}\tilde{U}\partial_{r}\tilde{U}+\partial_{r}P(\tilde{\rho})=(2\mu+\lambda)\partial_{r}\Big{(}\frac{(r^{n-1}\tilde{U})_{r}}{r^{n-1}}\Big{)}$ for $\text{for }\ r\geq 1,$ (1.7b)
$\tilde{U}(1)=u_{b},\quad\lim_{r\to\infty}\big{(}\tilde{\rho}(r),\tilde{U}(r)\big{)}=(\rho_{+},0).$ (1.7c)
Multiplying $r^{n-1}$ on (1.5a) and integrating the resultant equality over $(1,r)$, we obtain
@@
For general spatial dimension $n\geq 2$, I. Hashimoto and A. Matsumura in [3] proved the existence and uniqueness of a spherically symmetric classical solution to (1.5), provided that $|u_{b}|$ is small. Subsequently in [4], I. Hashimoto, S. Nishibata and S. Sugizaki refined the spatial decay estimate for $(\tilde{\rho},\tilde{U})$, and they also prove the asymptotic stability of $(\tilde{\rho},\tilde{U})$ for small, spherically symmetric initial perturbations. In particular, the well-posedness theorem for the stationary solution can be summarised as follows:
@@
There exists $\delta=\delta(\rho_{+},\mu,\gamma,K,n)>0$ such that if $|u_{b}|\leq\delta$, then the unique solution $(\tilde{\rho},\tilde{U})(r)\in\mathcal{C}^{2}[1,\infty)$ to the problem (1.7) exists. In addition, $\tilde{\rho}(r)$ and $\tilde{U}(r)$ are respectively strictly monotone increasing and decreasing. Moreover, there exists constant $C=C(\rho_{+},\mu,\gamma,K,n)>0$ such that for $r\geq 1$,
$\tilde{U}(r)=u_{b}\tilde{\rho}(1)\dfrac{r^{1-n}}{\tilde{\rho}(r)},\quad$ $|\tilde{\rho}(r)-\rho_{+}|\leq Cr^{-2n+2},$
$C^{-1}|u_{b}|r^{-n}\leq|\partial_{r}\tilde{U}(r)|\leq C|u_{b}|r^{-n},\quad$ $C^{-1}|u_{b}|^{2}r^{-2n+1}\leq|\partial_{r}\tilde{\rho}(r)|\leq C|u_{b}|^{2}r^{-2n+1},$
$C^{-1}|u_{b}|r^{-2n}\leq|\partial_{r}^{2}\tilde{U}(r)|\leq C|u_{b}|r^{-2n},\quad$ $C^{-1}|u_{b}|^{2}r^{-2n}\leq|\partial_{r}^{2}\tilde{\rho}(r)|\leq C|u_{b}|^{2}r^{-2n}.$
Y. Huang and S. Nishibata in [6] prove that if the initial data (1.3) is spherically symmetric, then the stationary solution $(\tilde{\rho},\tilde{u})$ obtained in Lemma 1.1 is still asymptotically stable even if the initial perturbation is large. However, the stability under general possibly non-spherically symmetric initial perturbation remained an open question. In this paper, we give a positive answer to this problem provided that the initial perturbation $(\rho_{0}-\tilde{\rho},u_{0}-\tilde{u})$ is small in some suitable Sobolev space.
2 Main Theorem
@@
then the initial boundary value problem (1.1)–(1.4) has a unique solution $(\rho,u)$ globally-in-time, which satisfies
@@
In addition, the following convergence holds,
$\lim\limits_{t\to\infty}\sup\limits_{x\in\Omega}|(\rho-\tilde{\rho},u-\tilde{u})(t,x)|=0.$
Remark 2.2.
The out-flowing boundary condition (1.2) plays a significant role in the derivation for $L^{\infty}_{T}H^{2}$-estimates of $\big{(}\nabla(\rho-\tilde{\rho}),\nabla(u-\tilde{u})\big{)}$. Since $\partial\Omega$ is a sphere, we observe that the temporal and angular derivatives preserve the boundary condition (1.2), while the radial derivative of $u$ at $\partial\Omega$ is not a-priori given. Due to this, it is necessary to split the analysis into 3 steps as follows: 1. obtain estimates for temporal and angular derivatives of $(\rho-\tilde{\rho},u-\tilde{u})$ using the parabolic structure of (1.1b); 2. obtain estimates for radial derivative of $\rho-\tilde{\rho}$ using the damping effect of compressible Navier-Stokes equations; 3. obtain estimates for full spatial derivatives of $(\rho-\tilde{\rho},u-\tilde{u})$ by treating the system (1.1) as Stokes equations. It is worth noting that in step 2 (Section 3.4), a cancellation occurs on the radial derivatives of $u$. This serves as the key argument to successfully close the estimate.
@@
The compressible Navier-Stokes equation is an important subject in mathematical physics. Particularly, the problems of out-flowing or in-flowing boundary have gained traction in recent decades. We state several previous results, which are relevant to the present paper.
@@
First, for a comprehensive survey of the mathematical theory of compressible Navier-Stokes equation, we refer readers to the book [7] by S. N. Antontsev, A. V. Kazhikhov, and V. N. Monakhov. The first breakthrough research on the large time stability of solution is done by A. Matsumura and T. Nishida in [15], where they consider the heat-conductive compressible flow for a general $3$-dimensional exterior domain with adhesion boundary condition ($u|_{\partial\Omega}=0$ and $\partial_{n}\theta|_{\partial\Omega}=0$ where $\theta$ is the absolute temperature). In this case the stationary solution is $(\tilde{\rho},0,\bar{\theta})$ where $\tilde{\rho}=\tilde{\rho}(x)$ is a positive function of spatial variable, and $\bar{\theta}$ is a positive constant. Under smallness assumptions on the initial perturbation and external forces, it is proved that $(\tilde{\rho},0,\bar{\theta})$ is stable globally in time.
@@
When the equation is spherically symmetric with adhesion boundary condition, a pioneering work has been done by N. Itaya [9], which establishes the global-in-time existence of a unique classical solution on a bounded annulus domain. In this work, no smallness assumption on the initial data is imposed. Later, T. Nagasawa studies the asymptotic state for the same problem in [20]. The paper [9] has motivated a sequence of developments on the topic of spherically symmetric solution. For instance, A. Matsumura in [12] constructed a spherically symmetric classical solution to the isothermal model with external forces on a bounded annular domain. Moreover, he also shows the convergence rate to the stationary solution as time tends to infinity is exponential. Subsequently, this result has been extended to the isentropic and heat-conductive models by K. Higuchi in [8]. The well-posedness of spherically symmetric solution in an unbounded exterior domain is first obtained by S. Jiang [10], where the global-in-time existence of a uniqueness classical solution is shown. In addition, a partial result on the time asymptotic stability is proved in [10] where, for $n=3$, $\|u(\cdot,t)\|_{L^{2j}}\to 0$ as $t\to\infty$ with any fixed integer $j\geq 2$. Later, this restriction on the long time stability was fully resolved by T. Nakamura and S. Nishibata in [18], where a complete stability theorem was obtained for both $n=2$ and $3$ with large initial data. We also refer to the paper [17] by T. Nakamura, S. Nishibata, and S. Yanagi, where the time asymptotic stability of the spherically symmetric solution for the isentropic flow is established with large initial data and external forces.
@@
For the general outflow or inflow problem, the stationary solution becomes non-trivial. This leads to a variety of physically interesting time-asymptotic behaviours for the solutions. A. Matsumura in [13] starts the first investigation of this problem for the isentropic model posed on the one dimensional half-space domain. Several kinds of boundary conditions were considered in [13], which includes inflow, outflow, and adhesion boundary conditions. He formulated conjectures on the classification of asymptotic behaviours of the solutions in different cases subject to the relation between the boundary data and the spatial asymptotic data. Then the stability theorems for some cases of inflow problem were established by A. Matsumura and K. Nishihara in [16] where they employed the Lagrangian mass coordinate. Following this work, S. Kawashima, S. Nishibata and P. Zhu in [11] further examine the outflow problem in one dimensional half space. They prove the long time stability of solutions with small initial perturbation with respect to the stationary solution. A detailed desicription of the convergence rate towards the stationary solution is found in the work of T. Nakamura, S. Nishibata and T. Yuge [17].
For the non-isentropic inflow problem in the half line, T. Nakamura and S. Nishibata in [19] demonstrate the time-asymptotic stability of stationary solutions under a small initial perturbation, for both the subsonic and transonic cases.
@@
The research on the outflow and inflow problems for the spherically symmetric solution in an unbounded exterior domain has been relatively new subject. Under the assumption that velocity at boundary, $|u_{b}|$ is sufficiently small, I. Hashimoto and A. Matsumura in [3] employ the iteration method to obtain the existence of a unique spherically symmetric stationary solutions for both inflow and outflow problems in an exterior domain. More recently, I. Hashimoto, S. Nishibata, S. Sugizaki in [4] show the stability of spherically symmetric stationary solution for outflow problem under small initial perturbation. Subsequently, Y. Huang and S. Nishibata in [6] considers the large, spherically symmetric initial perturbation for the outflow problem and they prove that the stationary solution remains stable in large time without any smallness assumption on the initial data. The main theorem proved in this paper is the first stability result for outflow problem under general possibly non-spherically symmetric initial perturbation.
@@
At the beginning of Section 3, we restate the equations (1.1) in terms of the difference functions $(\phi,\psi)=(\rho-\tilde{\rho},u-\tilde{u})$, and we reformulate them in linear forms which is suitable for the a-priori estimates. Moreover, we also provide the precise definition of energy norm $N(T)$ that is used throughout the proof. Then Section 3 is divided into 5 subsections as follows: 1. in Section 3.1, we derive the $L^{2}$–estimates of $(\phi,\psi)$ through the use of $L^{2}$ relative energy density, denoted $\mathcal{E}$; 2. in Sections 3.23.3, the estimates for temporal and angular derivatives are obtained using the parabolic structure of the momentum equation (1.1b); 3. in Section 3.4, we derive estimates involving the radial derivative of $\phi$, which is achieved via the coupling structure of continuity equation (1.1a) and momentum equation (1.1b). Finally in Section 3.5, we obtain estimates for the full derivatives of $(\phi,\psi)$, by treating (1.1b) as an elliptic system, and (1.1a)–(1.1b) as a Stokes-type system.
@@
    For $T\geq 0$, we denote the space-time domains as
@@
    For Cartesian coordinate $(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$, we denote $\hat{e}_{1}$, $\hat{e}_{2}$, $\hat{e}_{3}$ as the corresponding unit basis vectors. Moreover, for a scalar valued function $F\vcentcolon\mathbb{R}^{3}\to\mathbb{R}$ and a vector-valued function $V\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$ with $V=(V^{1},V^{2},V^{3})^{\top}$, we denote the differential operators:
    $\nabla F\vcentcolon=\sum_{i=1}^{3}\hat{e}_{i}\partial_{x_{i}}F,\quad$ $\textnormal{{div}}V\vcentcolon=\sum_{i=1}^{3}\partial_{x_{i}}V^{i},\quad$ $\Delta F\vcentcolon=\sum_{i=1}^{3}\partial_{x_{i}}^{2}F,$
    $(V\cdot\nabla)F\vcentcolon=\sum_{i=1}^{3}V^{i}\partial_{x_{i}}F,\quad$ $D_{t}F\vcentcolon=\partial_{t}F+(u\cdot\nabla)F.$
    In addition, for operators $S$, $T$ we denote $[S,T]\vcentcolon=ST-TS$ as the commutator.

### v2

@@ Abstract
[…]
@@ 1 Introduction
$\partial_{t}(\rho u)+\textnormal{{div}}(\rho u\otimes u)+\nabla P(\rho)=\mu\Delta u+(\mu+\lambda)\nabla\textnormal{{div}}u.$ (1.1b)
We study the asymptotic behavior of a solution $(\rho,u)$ to (1.1) in an unbounded exterior domain $\Omega\vcentcolon=\{x=(x_{1},x_{2},x_{3})^{\top}\in\mathbb{R}^{3}\;|\;|x|>1\}$. Here $\rho>0$ is the mass density; $u=(u^{1},u^{2},u^{3})^{\top}$ is the velocity of gas; $P(\rho)=K\rho^{\gamma}\;(K>0,\gamma\geq 1)$ is the pressure with the adiabatic exponent $\gamma$; $\mu$ and $\lambda$ are constants called viscosity-coefficients satisfying $\mu>0$ and $2\mu+3\lambda\geq 0$. Hereafter, we denote $Lu=\big((Lu)^{1},(Lu)^{2},(Lu)^{3}\big)^{\top}$ with
$(Lu)^{i}\vcentcolon=\mu\Delta u^{i}+(\mu+\lambda)\partial_{x_{i}}\textnormal{{div}}u=\sum_{j=1}^{3}\{\mu\partial_{x_{j}}^{2}u^{i}+(\mu+\lambda)\partial_{x_{i}}\partial_{x_{j}}u^{j}\}\ \text{ for }\ i=1,2,3.$
In the present paper we study the problem where fluid flows out at a constant velocity from an inner sphere of which center is at the origin with radius $|x|=1$. The corresponding boundary condition is given as
$\ \,u(t,x)=u_{b}\frac{x}{|x|}\ \$$\text{ for }\ x\in\partial\Omega,\$$\text{ where }\ u_{b}<0,$ (1.2a)
$\lim\limits_{|x|\to\infty}(\rho,u)(t,x)=(\rho_{+},0)\ \$$\text{ for }\ t\geq 0,\$$\text{ where }\ \rho_{+}>0.$ (1.2b)
For equations (1.1)–(1.2), we prescribe the initial data $(\rho_{0},u_{0})$ as
$\rho(x,0)=\rho_{0}(x)>0,\qquad u(x,0)=u_{0}(x),$ (1.3a)
$\lim_{|x|\to\infty}(\rho_{0}(x),u_{0}(x))=(\rho_{+},0),\qquad 0<\rho_{+}<\infty.$ (1.3b)
In addition, the data (1.3a) is assumed to be compatible with the boundary data (1.2a). Namely, we assume that
@@
Since the characteristic velocity of (1.1a) is directed towards the origin at the boundary due to (1.2a), no boundary condition for $\rho$ at $x\in\partial\Omega$ is necessary for the well-posedness of the initial boundary value problem (1.1)–(1.4).
1.1 Spehrically symmetric stationary solution
[…]
$\textnormal{{div}}\big(\tilde{\rho}\tilde{u}\big)=0$ $\text{for }\ x\in\Omega,$ (1.5a)
$\tilde{\rho}(\tilde{u}\cdot\nabla)\tilde{u}+\nabla P(\tilde{\rho})=L\tilde{u}$ $\text{for }\ x\in\Omega,$ (1.5b)
$\tilde{u}(x)=u_{b}\frac{x}{|x|}$ $\text{for }\ x\in\partial\Omega,$ (1.5c)
$(\tilde{\rho},\tilde{u})(x)\to(\rho_{+},0)$ $\text{as }\ |x|\to\infty.$ (1.5d)
[…]
$\tilde{\rho}(x)=\tilde{\rho}(r),\quad\tilde{u}(x)=\tilde{U}(r)\frac{x}{r},\qquad\text{where }\ r\vcentcolon=|x|\geq 1,$ (1.6)
for a certain function $\tilde{U}\vcentcolon[1,\infty)\to\mathbb{R}$. Then the system (1.5) is reformulated as
$\partial_{r}\big(r^{n-1}\tilde{\rho}\tilde{U}\big)=0$ $\text{for }\ r\geq 1,$ (1.7a)
$\tilde{\rho}\tilde{U}\partial_{r}\tilde{U}+\partial_{r}P(\tilde{\rho})=(2\mu+\lambda)\partial_{r}\Big(\frac{(r^{n-1}\tilde{U})_{r}}{r^{n-1}}\Big)$ $\text{for }\ r\geq 1,$ (1.7b)
$\tilde{U}(1)=u_{b},\qquad\lim_{r\to\infty}\big(\tilde{\rho}(r),\tilde{U}(r)\big)=(\rho_{+},0).$ (1.7c)
Multiplying the equation (1.5a) by $r^{n-1}$, and integrating the resultant equality over $(1,r)$, we obtain
@@ 1.1 Spehrically symmetric stationary solution
[…]
@@
There exists $\delta=\delta(\rho_{+},\mu,\gamma,K,n)>0$ such that if $|u_{b}|\leq\delta$, then the unique solution $(\tilde{\rho},\tilde{U})(r)\in\mathcal{C}^{2}[1,\infty)$ to the problem (1.7) exists. In addition, $\tilde{\rho}(r)$ and $\tilde{U}(r)$ are respectively strictly monotone increasing and decreasing. Moreover, there exists a constant $C=C(\rho_{+},\mu,\gamma,K,n)>0$ such that for $r\geq 1$,
$\tilde{U}(r)=u_{b}\tilde{\rho}(1)\dfrac{r^{1-n}}{\tilde{\rho}(r)},\qquad$ $|\tilde{\rho}(r)-\rho_{+}|\leq Cr^{-2n+2},$
$C^{-1}|u_{b}|r^{-n}\leq|\partial_{r}\tilde{U}(r)|\leq C|u_{b}|r^{-n},\qquad$ $C^{-1}|u_{b}|^{2}r^{-2n+1}\leq|\partial_{r}\tilde{\rho}(r)|\leq C|u_{b}|^{2}r^{-2n+1},$
$C^{-1}|u_{b}|r^{-n-1}\leq|\partial_{r}^{2}\tilde{U}(r)|\leq C|u_{b}|r^{-n-1},\qquad$ $C^{-1}|u_{b}|^{2}r^{-2n}\leq|\partial_{r}^{2}\tilde{\rho}(r)|\leq C|u_{b}|^{2}r^{-2n}.$
[…]
1.2 Main theorem
@@
then the initial boundary value problem (1.1)–(1.4) has a unique solution $(\rho,u)$ globally-in-time in the function space
@@
[…]
$\lim\limits_{t\to\infty}\sup\limits_{x\in\Omega}|(\rho-\tilde{\rho},u-\tilde{u})(t,x)|=0.$ (1.8)
@@
The well-posedness problem of compressible Navier-Stokes equation with out-flowing or in-flowing boundary conditions have gained traction in recent decades. We state several previous results which are relevant to the present paper.
@@
First, we refer readers to the book [2] by S. N. Antontsev, A. V. Kazhikhov, and V. N. Monakhov, where a comprehensive survey of the mathematical theory of compressible Navier-Stokes equation is given. For the heat-conducing model, A. Tani in [27] provides a comprehensive proof for the existence and uniqueness of a local-in-time classical solution in the Hölder sapce. The major breakthrough research on the time-asymptotic stability of solution is first accomplished by A. Matsumura and T. Nishida in [19], where they consider the heat-conductive compressible flow for a general $3$-dimensional exterior domain with external force and adhesion boundary condition ($u|_{\partial\Omega}=0$ and $\partial_{n}\theta|_{\partial\Omega}=0$ where $\theta$ is the absolute temperature). In this case, the stationary solution takes the form $(\tilde{\rho},0,\bar{\theta})$ where $\tilde{\rho}=\tilde{\rho}(x)$ is a positive function of spatial variable, and $\bar{\theta}$ is a positive constant. Under smallness assumptions on the initial perturbation, it is proved that $(\tilde{\rho},0,\bar{\theta})$ is stable globally in time.
@@
For spherically symmetric flow in the bounded annulus domains with adhesion boundary condition, a pioneering work has been done by N. Itaya [11], which establishes the global-in-time existence of a unique classical solution for the isothermal model. In this work, no smallness assumption on the initial data is imposed. Later, T. Nagasawa investigates the asymptotic stability for the same problem in [26]. The paper [11] has motivated a sequence of developments on the topic of spherically symmetric solution. For instance, A. Matsumura in [15] constructs a spherically symmetric classical solution to the isothermal model with external forces on a bounded annular domain. Moreover, he also proves that the convergence rate for the time-asymptotic stability is exponential. Subsequently, these result is extended to the isentropic and heat-conductive models by K. Higuchi in [10]. The well-posedness of a spherically symmetric solution in an unbounded exterior domain is first obtained by S. Jiang [12], where the global-in-time existence of the uniqueness classical solution is shown. In addition, a partial result on the time asymptotic stability is proved in [12] where, for $n=3$, $\|u(\cdot,t)\|_{L^{2j}}\to 0$ as $t\to\infty$ with an arbitrary fixed integer $j\geq 2$. Later, this restriction on the long time stability was fully resolved by T. Nakamura and S. Nishibata in [23], where a complete stability theorem is obtained for large initial data. We also refer to the paper [21] by T. Nakamura, S. Nishibata and S. Yanagi, where the time asymptotic stability of the spherically symmetric solution for the isentropic flow is established with large initial data and external forces.
@@
For the one-dimensional outflow or inflow problem, the stationary solution becomes non-trivial. This leads to a variety of physically interesting time-asymptotic behaviours for the solutions. A. Matsumura in [16] starts the first investigation of this problem for the isentropic model posed on the one dimensional half-space domain. Several kinds of boundary conditions are considered in [16], which includes inflow, outflow and adhesion boundary conditions. He formulates conjectures on the classification of asymptotic behaviours of the solutions in different cases subject to the relation between the boundary data and the spatial asymptotic data. Then the stability theorems for several cases of inflow problem were established by A. Matsumura and K. Nishihara in [20], where they employ the Lagrangian mass coordinate. Following this work, S. Kawashima, S. Nishibata and P. Zhu in [14] further study the outflow problem in one dimensional half space. They prove the long time stability of solutions with small initial perturbation from the stationary solution. A detailed desicription of the convergence rate towards the stationary solution is found in the work of T. Nakamura, S. Nishibata and T. Yuge [22]. For the non-isentropic inflow problem in the half line, T. Nakamura and S. Nishibata in [25] demonstrate the time-asymptotic stability of stationary solutions under a small initial perturbation, for both the subsonic and transonic cases.
@@
The research on the outflow and inflow problems for fluid flow in multi-dimensional (M-D for short) domains are relatively new subjects. For isentropic fluid occupying the M-D half space $\mathbb{R}^{n}_{+}$, Y. Kagei and S. Kawashima [13] show that the planar stationary solution for the out-flowing boundary condition is asymptotically stable. Subsequently T. Nakamura and S. Nishibata in [24] obtain the convergence rates for this time-asymptotic stability. In particular, they show that an algebraic or an exponential decay rate holds for the supersonic flow. Precisely, an algebraic convergence rate holds for the transonic flow provided that the respective initial perturbations decay in a normal direction to the boundary surface with the corresponding decay rates. For isentropic fluid occupying the exterior domain of a unit ball, I. Hashimoto and A. Matsumura in [5] obtain the existence of the unique spherically symmetric stationary solutions for both inflow and outflow problems provided that the velocity at boundary, $|u_{b}|$ is sufficiently small. More recently, I. Hashimoto, Y. Huang and S. Nishibata in [7] study the spherically symmetric inflow problem in the exterior domain and they prove that if the far-field state of the fluid is not vacuum but suitably rarefied, then the corresponding stationary solution is stable under small initial perturbation. For the exterior outflow problem, I. Hashimoto, S. Nishibata and S. Sugizaki in [6] show the stability of stationary solution with small spherically symmetric initial perturbations. Subsequently, this smallness assumption on initial data is removed by Y. Huang and S. Nishibata in [9] and they prove that the stationary solution for the outflow problem remains asymptotically stable under large spherically symmetric initial perturbations. In addition, E. Feireisl, P. Gwiazda, and A. Swierczewska-Gwiazda consider the finite energy weak solution to the exterior outflow problem in [3]. The main theorem proved in the present paper is the asymptotic stability for the outflow problem under general possibly non-spherically symmetric initial perturbations in the exterior domain.
@@
[…]
1. derive estimates for temporal and angular derivatives of $(\phi,\psi)$ by using the parabolic structure of (1.1b), which are the main contents of Sections 2.32.4;
2. derive estimates for radial derivative of $\phi$ using the damping effect of compressible Navier-Stokes equations, which is illustrated in Section 2.5;
3. derive estimates for the full spatial derivatives of $(\phi,\psi)$ by treating the system (1.1) as Stokes equations and this is devoted to Section 2.6. It is worth noting that in step 2 (Section 2.5), a cancellation occurs on the radial derivatives of $\psi$. This serves as the key argument to successfully close the estimate.
@@
    For $T\geq 0$, denote the space-time domains as
@@
    For the Cartesian coordinate $(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$, denote $\hat{e}_{1}$, $\hat{e}_{2}$, $\hat{e}_{3}$ as the corresponding unit basis vectors. Moreover, for functions $F\vcentcolon\mathbb{R}^{3}\to\mathbb{R}$, and $V\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$ with $V=(V^{1},V^{2},V^{3})^{\top}$, denote the differential operators:
    $\nabla F\vcentcolon=\sum_{i=1}^{3}\hat{e}_{i}\partial_{x_{i}}F,\qquad$ $\textnormal{{div}}V\vcentcolon=\sum_{i=1}^{3}\partial_{x_{i}}V^{i},\qquad$ $\Delta F\vcentcolon=\sum_{i=1}^{3}\partial_{x_{i}}^{2}F,$
    $(V\cdot\nabla)F\vcentcolon=\sum_{i=1}^{3}V^{i}\partial_{x_{i}}F,\qquad$ $D_{t}F\vcentcolon=\partial_{t}F+(u\cdot\nabla)F.$
    In addition, for operators $S$, $T$, denote $S,T]\vcentcolon=ST-TS$ as the commutator.
@@
    $H^{k}(\Omega)\!=\!\bigg\{F\in L^{2}(\Omega)\,\bigg|\,D^{k}F\vcentcolon=\Big\{\frac{\partial^{\alpha}F}{\partial x_{1}^{\alpha_{1}}\partial x_{2}^{\alpha_{2}}\partial x_{3}^{\alpha_{3}}},\,|\alpha|=l\Big\}\in L^{2}(\Omega),\ 1\leq l\leq k\bigg\},$
    Furthermore, the norms are defined by
    $\|F\|_{k}\vcentcolon=\|F\|_{H^{k}(\Omega)},\hskip 18.49988pt\|F\|\vcentcolon=\|F\|_{0}=\|F\|_{L^{2}(\Omega)},\hskip 18.49988pt\|F\|_{\infty}\vcentcolon=\operatorname*{\text{ess}\,\text{sup}}\limits_{x\in\Omega}|F|.$
2 A-priori estimates
2.1 Reformulation and energy norm $N(T)$
In this subsection, we reformulate the problem (1.1)-(1.3) to a form which is suitable for the a-priori estimate. In addition, we also define the notion of energy norm $N(T)$.